Solve this equation and show the work steps.
.5(x-9) = 3/4

Answer by donnieb82(2)

.5(x-9) = 3/4
0.5x-4.5 = 3/4 (Distribute)
4(0.5x-4.5)=4(3/4) (Multiply both sides by LCD (4) to clear any fractions)
2x-18=3 (Distribute & Multiply)
2x=3+18 (Add 18 to both sides)
2x=21 (Combine like terms on right side)
x=21/2 (Divide both sides by 2 to isolate x)
-------------
Answer or Solution is: x=21/2